Please help me solve this question! I have to have it solved by tomorrow for school or i will receive a failing grade for not completing it and getting the right answer. Thanks in advance for your help.
Find the zeros of the problem: x^3 - 10x^2 + 25x = 0
I came out with the solution (0,5,5) but i think it might actually be (0,5,D.R.)
*(D.R. meaning double root)(^ meaning to the __ power)

Answer by jim_thompson5910(35256)

x^3 - 10x^2 + 25x = 0

x(x^2 - 10x + 25) = 0

x(x - 5)(x - 5) = 0

x = 0 or x-5 = 0 or x-5 = 0

x = 0 or x = 5 or x = 5

So the solutions are x = 0 or x = 5 or x = 5

You can say that the unique roots are 0 and 5, where 0 is a single root and 5 is a double root.